# Metric Rarity and the Rise of Symmetry in Invariant Optimization

### v1

@@ Abstract
In the non-convex landscapes of $G$-invariant optimization problems—spanning particle physics, neural networks, and algebraic models—symmetric critical points defy statistical intuition. While generic configurations dominate the space, empirical evidence reveals two striking regimes: (I) symmetry as the norm, with asymmetric points vanishingly rare, and (II) a profound energetic ordering, where global minima exhibit disproportionately higher symmetry.
This paper unveils a unified geometric mechanism rooted in algebraic quotients. We reframe the problem on the quotient space $Y=X\mathbin{/\mkern-6.0mu/}G$, where the physical domain is the real image $L=\pi(X(\mathbb{R}))$—a metrically rare subset within the ambient quotient $Y(\mathbb{R})$. Theorems quantify this rarity: for the symmetric group $S_{n}$, the volume of the real image $L$ decays exponentially as $e^{-Cn\log n}$.
Regime I emerges from an “empty interior”: critical points evade the smooth locus of $L$, settling on boundaries tied to non-trivial stabilizers. Regime II invokes the “Active Constraint”—a global gradient channeling minima to sharp corners of $L$, manifesting as funnel topographies in physical systems that intercept the descent at crystalline structures with high symmetry.
Keywords: machine learning; symmetry breaking; invariant optimization; nonconvex loss landscapes; permutation symmetry; orbit space; algebraic quotients; stabilizer (isotropy) groups; real locus / real image; Lennard–Jones clusters.
@@ 1.1 Problem Statement: The Prevalence of Symmetry
Let $X$ be an irreducible complex affine algebraic variety defined over $\mathbb{R}$, equipped with a faithful action of a finite group $G$ by real automorphisms. We denote by $X(\mathbb{R})$ the set of real points of $X$; throughout this paper, we will refer to elements of $X$ as configurations. For any configuration $x\in X$, let $G_{x}=\{g\in G\mid g\cdot x=x\}$ denote its stabilizer subgroup. We classify a configuration $x$ as symmetric if its stabilizer $G_{x}$ is non-trivial, and asymmetric (or in general position) otherwise.
@@
Let $f:X\to\mathbb{C}$ be a $G$-invariant regular function defined over $\mathbb{R}$ (i.e., $f\in\mathbb{R}[X]^{G}$). Throughout this paper, we will denote the restriction of $f$ to the real locus $X(\mathbb{R})$ simply by $f$, distinguishing between the complex and real functions only when necessary for clarity. Our primary focus is the behavior of $f$ on the real locus $X(\mathbb{R})$. Within this setting, a pivotal question regarding the critical points is the emergence of symmetry: to what extent do these points possess non-trivial stabilizers?
@@
The existence of symmetric critical points is, in itself, mathematically unremarkable. By the Principle of Symmetric Criticality (see Theorem 2.6), any critical point of $f$ restricted to a real fixed-point subspace (definition 2.3) is necessarily a critical point of $f$ on the entire real manifold; thus, the existence of symmetric critical points is almost guaranteed.
@@
Consider the geometry of the real locus $X(\mathbb{R})$. Within this manifold, the set of symmetric configurations forms a lower-dimensional subvariety of measure zero. Conversely, the asymmetric configurations (those in general position) constitute an open, dense set, occupying the full measure of the configuration space. Now, consider a rugged $G$ invariant energy landscape $f:X(\mathbb{R})\to\mathbb{R}$ characterized by a proliferation of critical points. Statistically, one would expect the overwhelming majority of these points to inhabit the generic, asymmetric stratum. Moreover, this statistical dominance implies a specific prediction for the ground state: assuming the energy values are distributed randomly across the critical points, the probability that the global minimum coincides with a rare symmetric configuration should be negligible.
@@
Empirical observation, however, defies this expectation on two levels.
@@
First, in what we term Regime I, symmetry appears to be ubiquitous rather than rare. This phenomenon was rigorously documented by Arjevani et al. [1, 2], who observed that loss functions of shallow ReLU networks systematically preserve the permutation symmetries of hidden neurons, and that large isotropy groups similarly persist in the landscape of symmetric tensor decomposition.
@@
In companion numerical experiments [3], we investigated this question across a broad family of invariant optimization problems of the following common form: a finite group $G$ acts faithfully and linearly on a real vector space $X(\mathbb{R})$, and $f:X(\mathbb{R})\to\mathbb{R}$ is a $G$-invariant polynomial. This framework simultaneously covers, for example, particle interaction models (with $G=S_{n}$ acting by label permutations), graph-based objective functions with automorphism symmetries, and invariant functions on finite projective geometries (with $G$ a finite linear group such as $PGL_{n}(\mathbb{F}_{q})$ acting on the relevant representation space). Across these algebraically distinct settings— and especially at higher polynomial degrees —we observed that almost all numerically detected critical points had non-trivial stabilizers, in sharp contrast with the naive measure-theoretic heuristic that symmetric configurations should be negligible.
@@
Second, and perhaps more fundamentally, distinct preference for higher symmetry emerges at the energetic extremes—a phenomenon we term Regime II. Even in the pervasive cases mentioned above where symmetry is the norm, the deepest critical points exhibit the largest stabilizers (Figure 1(a)). Additional numerical evidence and robustness checks appear in Appendix B.
@@
Observation: A stark phase separation is visible. The deep energy funnel ($t\to 0$) exhibits a high concentration of symmetric configurations (up to 20% density). As energy increases, the probability of encountering symmetry decays rapidly towards the global baseline (dashed red line), confirming that symmetry is structurally linked to low-energy order rather than being uniformly distributed.
@@
This correlation becomes the decisive selection principle in complex physical systems like Lennard-Jones (LJ) clusters, governed by the potential $V_{LJ}$ (Definition 5.2). These systems map naturally to our algebraic framework: the configuration space modulo global motions acts as the variety $X_{\operatorname{shape}}$ (Definition 5.4), and the physical point group is identified with the stabilizer in $S_{n}$ (Proposition 5.6).
In this rugged setting, the asymmetric “ocean” is vast: for 13 particles (LJ13), approximately 94% of the 30,517 known critical points and approximately 78% of the local minima documented in the Cambridge Cluster Database [^1] https://www-wales.ch.cam.ac.uk/CCD.html possess a trivial stabilizer. Thus, the first phenomenon (ubiquity) is absent. However, the drive toward symmetry at the extremes returns decisively. As illustrated in Figure 1(b), the local density of symmetric configurations rises dramatically as one moves deeper into the potential wells, culminating in the global minimum: the regular icosahedron (order 120). This prevalence of symmetry is not unique to $N=13$; the pattern persists across the range $LJ_{1}$$LJ_{150}$, where approximately 83% of global minima possess non-trivial point-group symmetry [4].
The systematic correlation between high symmetry and energetic extremes—specifically the tendency of symmetric structures to occupy the tails of the energy distribution—was originally documented by Wales [5]. Through extensive surveys of model systems, Wales established that while symmetry might not dictate the average energy, it is a defining characteristic of the lowest and highest energy states. It is important to distinguish the scopes of these inquiries: while Wales analyzed near-symmetry (defining symmetry via continuous measures and tolerance thresholds), our framework focuses on exact symmetry. Remarkably, the preference for symmetry at energetic extremes remains robust even under this strict algebraic definition.
@@ 1.2 Our Proposal: Geometry of the Real Image
We shift our analysis from the configuration space $X$ to the quotient space $Y=X\mathbin{/\mkern-6.0mu/}G$. Let $\pi:X\to Y$ denote the quotient map (see Section 2). The physical configurations form the real image $L=\pi(X(\mathbb{R}))$, a semi-algebraic subset of the real quotient $Y(\mathbb{R})$. There is a canonical isomorphism between $G$-invariant functions on $X$ and regular functions on $Y$. Any invariant energy $f$ factors as $f=\tilde{f}\circ\pi$ for a unique $\tilde{f}:Y(\mathbb{R})\to\mathbb{R}$. Crucially, this implies that invariance imposes no structural constraint on $\tilde{f}$ itself; $\tilde{f}$ can be any generic function. Consequently, the landscape features of $f$ are determined solely by the interaction between the function $\tilde{f}$ and the geometry of the domain $L$.
@@
Let $x$ be a smooth point of $X$. We prove in Theorem 3.1 that the differential $d\pi_{x}$ is surjective if and only if the stabilizer $G_{x}$ is trivial. This implies a fundamental distinction:
@@
We prove in Theorem 3.2 that the topology of $L$ is linked to symmetry: points in the interior $L^{\circ}$ correspond to odd-order stabilizers, while points on the boundary $\partial L$ correspond to even-order stabilizers. In reflection groups such as $S_{n}$, every non-trivial stabilizer has even order, therefore $\pi$ maps all symmetric points to the boundary $\partial L$. A similar restriction appears empirically for Lennard–Jones critical points, where odd-order symmetries are exceptionally rare (a phenomenon we discuss later). Consequently, throughout this paper we will adopt a convention of brevity: we will frequently describe the properties of a configuration $x$ directly in terms of its image $\pi(x)$, referring to symmetric configurations simply as lying on the boundary $\partial L$.
The question becomes: Why do critical points exhibit a statistical preference for the boundary $\partial L$? This preference manifests in varying degrees: from cases where the interior is effectively empty (Regime I), to cases where the boundary simply captures the global minimum (Regime II).
@@ 1.2.1 The Metric Rarity of the Real Image
The dichotomy between the “Smooth Region” (trivial stabilizer, interior $L^{\circ}$) and the “Singular Region” (non-trivial stabilizer, boundary $\partial L$) derives its significance from a fundamental geometric reality: the real image $L$ is metrically rare within the quotient space $Y(\mathbb{R})$. While $L$ and $Y(\mathbb{R})$ share the same topological dimension, the volume of $L=\pi(X(\mathbb{R}))$ becomes vanishingly small as the system size increases. This rarity is universal across our models.
@@
Now, let us define a Gaussian measure on the quotient space $Y(\mathbb{R})\cong\mathbb{R}^{n}$. Under the hypothesis that the Monic and Kac Gaussian ensembles share the same asymptotic decay rate (Assumption 4.1), Theorem 4.2 implies that the measure of the real image $L$ decays super-exponentially ($\sim e^{-Cn^{2}}$).
@@
Crucially, this rarity extends to the setting of interacting particles. In this context, we formulate the system over the physical shape space $X_{\operatorname{shape}}^{+}\subset X_{\operatorname{shape}}(\mathbb{R})$, defined by particle positions modulo global motions (see Definition 5.4). The geometric structure is established by assuming any Hermitian metric on the ambient configuration space of centered particle systems $\mathcal{V}=\mathbb{C}^{d\times(n-1)}$. This metric descends canonically to the Root-Mean-Square Deviation (RMSD) metric on the quotient $X_{\operatorname{shape}}^{\mathrm{\operatorname{pr}}}$ (see Definition 5.7), which subsequently induces a natural measure $\mu$ on the quotient space $Y_{\operatorname{shape}}(\mathbb{R})$ (as constructed in Definition 5.9).
@@
1.2.2 Regime I: The Empty Interior (Ubiquitous Symmetry)
In polynomial optimization problems invariant under the action of a finite group, as well as in shallow neural networks invariant under permutations of hidden neurons (Arjevani et al. [1, 2], Schneider [3] and Figure 1(a)), we observe ubiquitous symmetry. Geometrically, this corresponds to the scenario where the descended function $\tilde{f}$ possesses no critical points within the smooth interior of the real image $L$. We propose that this arises from a fundamental “size mismatch”: since the physical domain $L$ occupies a vanishingly small volume relative to the ambient quotient, the strict condition $d\tilde{f}=0$ is statistically unlikely to be satisfied within the interior $L^{\circ}$. Consequently, the critical points of $f$ arise exclusively from the singularities of the quotient map $\pi$—which correspond precisely to configurations with non-trivial stabilizers. We demonstrate this mechanism in the following heuristic example: let $f:\mathbb{R}^{n}\to\mathbb{R}$ be an $S_{n}$-invariant polynomial and assume that $\tilde{f}$ of degree $n$, then assume that $\tilde{f}$ has $n^{n}$ critical points (the generic upper bound for polynomials of degree $n$ in $n$ variables). Now assume that those critical points are ”normally” distributed within $Y(\mathbb{R})$. In Corollary 4.3, we show that the probability that there will be any critical point of $\tilde{f}$ in the real image $L$ decays to zero as $n$ increases. Therefore, if one accepts this normality assumption, it is not surprising that for large $n$, there is no critical point of $f$ with a trivial stabilizer.

### v2

# Metric Rarity and the Emergence of Symmetry in G-Invariant Potential Surfaces

@@ Abstract
Let $X$ be an irreducible complex affine algebraic variety defined over $\mathbb{R}$, equipped with a faithful action of a finite group $G$, and let $Y=X\mathbin{/\mkern-6.0mu/}G$ denote the categorical quotient with projection $\pi$. We study the geometry of the real image $L=\pi(X(\mathbb{R}))\subset Y(\mathbb{R})$ and its consequences for $G$-invariant optimization.
Equipping $Y(\mathbb{R})$ with the measure induced by a $G$-invariant metric on $X$, we prove that the relative volume of $L$ in $Y(\mathbb{R})$ equals $(\#\mathrm{Inv}(G))^{-1}$, where $\mathrm{Inv}(G)$ is the set of involutions of $G$. For the symmetric group $S_{n}$ acting on $\mathbb{R}^{n}$, this ratio decays super-exponentially in $n$. In particular, $L$ is metrically rare within the ambient real quotient.
We apply this result to two phenomena observed in $G$-invariant optimization problems:
Regime I (Rarity of asymmetric critical points). The super-exponential decay of the volume of $L$ renders the interior $L^{\circ}$ statistically negligible as a locus for critical points. This geometric rarity provides a rationale for the observed prevalence of symmetry: generic critical points are constrained to the boundary strata of $L$, corresponding to orbits with non-trivial stabilizers.
Regime II (Energetic ordering by symmetry). We formulate the Active Constraint hypothesis: due to the metric rarity of the real image $L$, the landscape is dominated by a global gradient that drives the deepest descent trajectories toward the boundary of $L$. This global gradient directs the global minimum into the high-codimension strata of the boundary—corresponding to large stabilizers—thereby establishing a structural link between low energy and non-trivial stabilizers. This mechanism rationalizes the funnel topography of Lennard-Jones clusters, where the system is funneled into a crystallized ground state.
@@ 1.1 Problem Statement: The Prevalence of Symmetry
Let $X$ be an irreducible complex affine algebraic variety defined over $\mathbb{R}$, equipped with a faithful action of a finite group $G$ by real automorphisms. We denote by $X(\mathbb{R})$ the set of real points of $X$; we refer to points $x\in X$ as configurations. For any configuration $x\in X$, let $G_{x}=\{g\in G\mid g\cdot x=x\}$ denote its stabilizer subgroup. We classify a configuration $x$ as symmetric if its stabilizer $G_{x}$ is non-trivial, and asymmetric (or in general position) otherwise.
@@
Let $f:X\to\mathbb{C}$ be a $G$-invariant regular function defined over $\mathbb{R}$ (i.e., $f\in\mathbb{R}[X]^{G}$). Throughout this paper, we will denote the restriction of $f$ to the real locus $X(\mathbb{R})$ simply by $f$, distinguishing between the complex and real functions only when necessary for clarity. Our primary focus is the behavior of $f$ on the real locus $X(\mathbb{R})$. Within this setting, a central question regarding the critical points is the emergence of symmetry: we investigate the extent to which these points possess non-trivial stabilizers.
@@
The existence of symmetric critical points is, in itself, a standard consequence of the group action. By the Principle of Symmetric Criticality (see Theorem 2.6), any critical point of $f$ restricted to a real fixed-point subspace (definition 2.3) is necessarily a critical point of $f$ on the entire real manifold; thus, the existence of symmetric critical points is almost guaranteed.
@@
Consider the geometry of the real locus $X(\mathbb{R})$. Within this manifold, the set of symmetric configurations forms a lower-dimensional subvariety of measure zero. Conversely, the asymmetric configurations (those in general position) constitute an open, dense set, occupying the full measure of the configuration space. Consider a $G$-invariant energy landscape $f:X(\mathbb{R})\to\mathbb{R}$ with a large number of critical points. Statistically, one would expect the majority of these points to lie in the generic, asymmetric stratum. Moreover, this statistical dominance implies a specific prediction for the ground state: assuming the energy values are distributed randomly across the critical points, the probability that the global minimum coincides with a rare symmetric configuration should be negligible.
@@
Empirical observation, however, reveals two distinct deviations from this expectation.
@@
First, in what we term Regime I, symmetry appears to be prevalent rather than rare. This phenomenon was rigorously documented by Arjevani et al. [1, 2], who observed in the landscapes of shallow ReLU networks systematically preserve the permutation symmetries of hidden neurons, and that large isotropy groups similarly persist in the landscape of symmetric tensor decomposition.
@@
In companion numerical experiments [3], we investigated this question across a broad class of $G$-invariant variational problems of the following common form: a finite group $G$ acts faithfully and linearly on a real vector space $X(\mathbb{R})$, and $f:X(\mathbb{R})\to\mathbb{R}$ is a $G$-invariant polynomial. This framework simultaneously covers, for example, particle interaction models (with $G=S_{n}$ acting by label permutations), graph-based objective functions with automorphism symmetries, and invariant functions on finite projective geometries (with $G$ a finite linear group such as $PGL_{n}(\mathbb{F}_{q})$ acting on the relevant representation space). Across these algebraically distinct settings— and especially at higher polynomial degrees —we observed that almost all numerically detected critical points had non-trivial stabilizers, contrary to the statistical expectation.
@@
Second, a distinct preference for higher symmetry emerges at the energetic extremes—a phenomenon we term Regime II. Even in the cases mentioned above where symmetry is the norm, the deepest critical points exhibit the largest stabilizers (Figure 1(a)). Additional numerical evidence and robustness checks appear in Appendix B.
@@
Observation: A clear separation between two regimes is visible. The deep energy funnel ($t\to 0$) exhibits a high concentration of symmetric configurations (up to 20% density). As energy increases, the probability of encountering symmetry decays rapidly towards the global baseline (dashed red line), confirming that symmetry is structurally linked to low-energy order rather than being uniformly distributed.
@@
This correlation becomes the dominant selection mechanism in complex physical systems such as Lennard-Jones (LJ) clusters, governed by the potential $V_{LJ}$ (Definition 5.2). These systems map naturally to our algebraic framework: the configuration space modulo global motions acts as the variety $X_{\operatorname{shape}}$ (Definition 5.4), and the physical point group is identified with the stabilizer in $S_{n}$ (Proposition 5.6).
In this setting, the generic stratum has full measure: for 13 particles (LJ13), approximately 94% of the 30,517 known critical points and approximately 78% of the local minima documented in the Cambridge Cluster Database [^1] https://www-wales.ch.cam.ac.uk/CCD.html possess a trivial stabilizer. Thus, the first phenomenon (prevalence) is absent. However, the drive toward symmetry at the extremes re-emerges. As illustrated in Figure 1(b), the local density of symmetric configurations increases monotonically as one moves deeper into the potential wells, culminating in the global minimum: the regular icosahedron (order 120). This prevalence of symmetry is not unique to $N=13$; the pattern persists across the range $LJ_{1}$$LJ_{150}$, where approximately 83% of global minima possess non-trivial point-group symmetry [4].
The systematic correlation between high symmetry and energetic extremes—specifically the tendency of symmetric structures to occupy the tails of the energy distribution—was originally documented by Wales [6]. Through extensive surveys of model systems, Wales established that while symmetry might not dictate the average energy, it is a defining characteristic of the lowest and highest energy states. We note that Wales analyzed near-symmetry (defining symmetry via continuous measures and tolerance thresholds), whereas our framework focuses on exact symmetry. Notably, the selection of symmetry at energetic extremes remains robust even under this strict algebraic definition.
@@ 1.2 Our Proposal: Geometry of the Real Image
We shift our analysis from the configuration space $X$ to the quotient space $Y=X\mathbin{/\mkern-6.0mu/}G$. Let $\pi:X\to Y$ denote the quotient map (see Section 2). The physical configurations form the real image $L=\pi(X(\mathbb{R}))$, a semi-algebraic subset of the real quotient $Y(\mathbb{R})$. There is a canonical isomorphism between $G$-invariant functions on $X$ and regular functions on $Y$. Any invariant energy $f$ factors as $f=\tilde{f}\circ\pi$ for a unique $\tilde{f}:Y(\mathbb{R})\to\mathbb{R}$. In particular, this implies that invariance imposes no structural constraint on $\tilde{f}$ itself; $\tilde{f}$ can be any generic function. Consequently, the landscape features of $f$ are determined solely by the interaction between the function $\tilde{f}$ and the geometry of the domain $L$.
@@
Let $x$ be a smooth point of $X$. Proposition 3.1 states that the differential $d\pi_{x}$ is surjective if and only if the stabilizer $G_{x}$ is trivial. This implies a fundamental distinction:
@@
Utilizing the stratification analysis of Procesi and Schwarz [7], we identify the topological boundary $\partial L$ with the locus of configurations possessing even-order stabilizers (Theorem 3.2). In reflection groups such as $S_{n}$, every non-trivial stabilizer has even order; consequently, $\pi$ maps all symmetric points to the boundary $\partial L$. A similar restriction appears empirically for Lennard–Jones critical points, where odd-order symmetries are exceptionally rare (a phenomenon we discuss later). Therefore, throughout this paper we adopt a convention of brevity: we frequently describe the properties of a configuration $x$ directly in terms of its image $\pi(x)$, referring to symmetric configurations simply as lying on the boundary $\partial L$.
We investigate the origin of this statistical concentration of critical points on the boundary $\partial L$. This tendency manifests in varying degrees: from cases where the interior $L^{\circ}$ is devoid of critical points (Regime I), to cases where the boundary simply captures the global minimum (Regime II).
@@ 1.2.1 The Metric Rarity of the Real Image
The dichotomy between the “Smooth Region” (trivial stabilizer, interior $L^{\circ}$) and the “Singular Region” (non-trivial stabilizer, boundary $\partial L$) derives its significance from the following geometric fact: the real image $L$ is metrically rare within the quotient space $Y(\mathbb{R})$. While $L$ and $Y(\mathbb{R})$ share the same topological dimension, the volume of $L=\pi(X(\mathbb{R}))$ becomes increasingly small as the system size grows. This rarity is universal across our models.
@@
We equip the quotient space $Y(\mathbb{R})\cong\mathbb{R}^{n}$ with a Gaussian measure. Under the hypothesis that the Monic and Kac Gaussian ensembles share the same asymptotic decay rate (Assumption 4.1), Theorem 4.2 implies that the measure of the real image $L$ decays super-exponentially ($\sim e^{-Cn^{2}}$).
@@
This rarity extends to the setting of interacting particles. In this context, we formulate the system over the physical shape space $X_{\operatorname{shape}}^{+}\subset X_{\operatorname{shape}}(\mathbb{R})$, defined by particle positions modulo global motions (see Definition 5.4). The geometric structure is established by assuming any Hermitian metric on the ambient configuration space of centered particle systems $\mathcal{V}=\mathbb{C}^{d\times(n-1)}$. This metric descends canonically to the Root-Mean-Square Deviation (RMSD) metric on the quotient $X_{\operatorname{shape}}^{\mathrm{\operatorname{pr}}}$ (see Definition 5.7), which subsequently induces a natural measure $\mu$ on the quotient space $Y_{\operatorname{shape}}(\mathbb{R})$ (as constructed in Definition 5.9).
@@
1.2.2 Regime I: The “Empty Interior” (No Interior Critical Points)
In polynomial optimization problems invariant under the action of a finite group, as well as in shallow neural networks invariant under permutations of hidden neurons (Arjevani et al. [1, 2], Schneider [3] and Figure 1(a)), we observe prevalent symmetry. Geometrically, this corresponds to the scenario where the descended function $\tilde{f}$ possesses no critical points within the smooth interior of the real image $L$. Equivalently, the condition $d\tilde{f}=0$ has no solutions in $L^{\circ}$. We propose that this arises from a volume disparity: since the physical domain $L$ occupies an increasingly small volume relative to the ambient quotient, the strict condition $d\tilde{f}=0$ is statistically unlikely to be satisfied within the interior $L^{\circ}$. Consequently, the critical points of $f$ arise exclusively from the singularities of the quotient map $\pi$—which correspond precisely to configurations with non-trivial stabilizers.
We illustrate this mechanism via a heuristic example. Consider an $S_{n}$-invariant polynomial $f:\mathbb{R}^{n}\to\mathbb{R}$ such that its quotient descendant $\tilde{f}$ is of degree $n$. By Bézout’s theorem, such a function possesses at most $n^{n}$ critical points. Proceeding under Assumption 4.1—specifically, that the real-root probability for the Monic Gaussian ensemble decays asymptotically as in the Kac ensemble—Corollary 4.3 establishes that the measure of the physical domain $L$ decays super-exponentially. Consequently, the expected number of critical points falling within the interior $L^{\circ}$ vanishes, implying a statistical prohibition on trivial stabilizers in the asymptotic limit.